OpenStudy (anonymous):

The length of a rectangle is 2 more than three times the width. The area of the rectangle is 161 square inches. What are the dimensions of the rectangle? If x = the width of the rectangle, which of the following equations is used in the process of solving this problem? 2x^2 + 3x - 161 = 0 3x^2 + 2x - 161 = 0 6x^2 - 161 = 0

OpenStudy (anonymous):

A=L*W L=2+3w 161=(2+3W)*(W) 161=2W+3W^2 3W^2+2W-161=0

OpenStudy (anonymous):

if I desever a medal then click good answer

OpenStudy (anonymous):

:) everyone deserves a medal :D